# MMLongBench: Benchmarking Long-Context Vision-Language Models Effectively and Thoroughly

### v1

@@ A.5 Long-Document VQA
First, we find that there is a proportion of noisy answer format annotations. For example, a list answer like [‘Top 10 File Categories Sorted By Disk Space’, ‘Last 12 Months Modified Disk Space History’] is annotated as being in string format, and answers in string format are also written in list format, such as [‘PRIVACY SCREEN OPTIONS’]. Also, percentages like 86% and 52% are labeled as float. Therefore, our first step with these datasets is to correct the mislabeled answer formats and discard the instances for which the correct answer format cannot be recovered.
Second, both datasets rely heavily on LLMs, such as GPT-4o, to extract the answer from model predictions. This leads to high evaluation costs and poses challenges for large-scale evaluation, like 46 models in our work. Then, we take a closer examination of different formats of answers: (1) For integer and float answers, we find that numbers can be extracted with regular expressions; (2) For string answers, if the answer is short (less than 5 words), we find that model predictions are also short. Thus, we can directly use automatic metrics like ROUGE without the need for answer extractions; (3) As a result, only long-form string answers require LLM-based extraction; Since long-form string answers ($>$ 5 words) constitute only a small proportion of these datasets, we simply discard those instances to enable scalable evaluation without relying on GPT-4o. We find that the retrained short string answers are mostly entity names; (4) Note that for list answers, we evaluate each element in the list (i.e., integer, float, or string), and the answer extraction method is determined by the type of each element.
@@
We follow previous works [5, 17] and employ the same rule-based scoring method that applies different strategies depending on the format of the reference answer: (1) For String format answers, we initially use regular expressions to determine whether the answers require exact matching (e.g., telephone numbers, email addresses, website addresses, filenames, times, dates, etc.). If the answer needs an exact match, we perform a substring exact matching (SubEM) with a score of 0 or 1. Otherwise, we follow previous works [81, 62, 96, 20] and calculate ROUGE F1 scores; (2) For integer answers, we perform an exact match comparison, and the score is either 0 or 1; (3) For float answers, we treat the model prediction and gold reference as the same if the relative error is less than 1%; (4) For list answers, we evaluate each element separately based on its answer type and take the average. Here, we follow LongDocURL to use the Greedy List Match: for each element in the reference list, we compute its score against each element in the prediction list and greedily select the highest score as its matching score. This metric does not require the predicted list to follow the same element order as the reference list, thereby providing greater tolerance in evaluation. After all the filtering, we find both datasets remain well-distributed as shown in Figure 9.
@@
The input lengths of DocVQA tasks are also easy to control. If an example exceeds a given length $L$, we truncate the document evenly from both sides while preserving the answer pages. If the document cannot fill the length $L$, we alternately pad the left and right sides with randomly sampled negative documents until the required length is reached. Notably, we also truncate the last padding document as needed to control the length at the granularity of pages, instead of documents.
@@ A.6 Image Resizing and Statistics
The number of tokens per image is determined by the image size in our benchmark, as we discussed in Section 3.3. In MM-NIAH, we find that many images from the OBELICS dataset are unnecessarily large (up to 8000$\times$6000 pixels) and are not text-rich. Then, we resize those images’ long edge to 1024 pixels while preserving their aspect ratio.
@@ B.1 LVLMs Beyond our Evaluation
DeepSeek-VL [102], DeepSeek-VL2 [48], and Long-Llava [53] are also excellent, but they present additional challenges. These models do not provide a standard API in the Transformers framework [103], which we used for inference. Therefore, these models cannot be loaded directly via Transformes, and we need to develop them based on their GitHub repositories. As a result, integrating these models requires a prohibitive amount of engineering effort to adapt their codebases, making it impractical within our current scope. We leave their integration for future work.
@@ Appendix C Experimental Setup
For all open-source models, we evaluate them on an 8$\times$×A100 (80GB each) GPU server. We use the HuggingFace framework [103] to deploy models and generate outputs. Since all models are instruction-tuned, we apply the chat templates to all datasets. We load models in BF16 with FlashAttention2 [104] for faster inference. The largest open-source models tested in our work have 72B parameters. Our computational resources are limited to 8 $\times$ A100 GPUs; thereby, we cannot evaluate models with over 100B parameters, such as Llama4 [4] at 128K tokens.
We sampled 100 examples from each dataset to evaluate models. This amount actually results in 600 examples for single-needle tasks: ViQuAE, VH-Single, and MM-NIAH-Ret, and 300 examples for multi-needle tasks: InfoSeek, VH-Multi, and MM-NIAH-Count. This is because we test 6 different depths (i.e., [0, 0.2, 0.4, 0.6, 0.8, 1.0]) for single-needle examples and 3 different permutations for multi-needle examples to mitigate the positional bias. Note that in MM-NIAH, we sample the text-needle and image-needle examples evenly, with 50 of each type. The MM-NIAH-Reason is more complex since image-needle (I) examples have a single needle, while text-needle (T) ones have multiple needles. There are 300 examples for MM-NIAH-Reason (I) (50 $\times$ 6 depths) and 150 examples for MM-NIAH-Reason (T) (50 $\times$ 3 permutations). Due to the depth imbalance in MM-NIAH-Reason, we compute the average score for each subset separately and report their mean as the final result.
Appendix D Addtional Results
@@ D.1 The Difficulty of Visual Haystack
We manually checked a number of examples from the Visual Haystack dataset and didn’t find any errors in the task labels. As shown in Figure 10, Gemini-2.5-Pro achieves an accuracy of 85.4 on VH-Single, and GPT-4o achieves an accuracy of 70.3 on VH-Multi. These results are significantly higher than a random guess (50%), which demonstrates the correctness of the dataset labels and our implementation.
@@
D.2 Correlation between NIAH and Variou Downstream Tasks
@@
We further show the performance of the selected models on the text-needle and image-needle subsets in Figure 12. From it, we observe that models exhibit weak performance on MM-NIAH-Count (I) and MM-NIAH-Reason (I). This challenging nature leads to a low degree of separability between different models.
@@ D.4 Performance of Claude
At the time of our evaluation, the Claude 3 family of models can take up to 100 images [^6] https://docs.anthropic.com/en/docs/build-with-claude/vision#basics-and-limits in a single request. However, a few datasets, such as Food101, VH-Single, or SlildeVQA, contain hundreds of images at input lengths of 64K and 128K tokens. As a result, it is impossible to process all images in a single pass through the model. For each input length, if one or more datasets within a category contain samples with more than 100 images, we exclude that category from evaluation at that input length. We provide the statistics about the average image number per example in each dataset at all five input lengths in Table 3 and Section A.6.
When testing Claude-3.7-Sonnet, we mark those untestable cases as “N/A” in our results to distinguish them from genuine model failures (which receive a score of 0).
Figure 14: Results of applying YaRN [37] to Qwen2.5-VL models. We find that YaRN substantially improves the performance of Qwen2.5-VL-3B. However, the SoTA performance from larger models (i.e., 32B and 72B) only fluctuates slightly.
@@ D.5 Positional Embedding Extrapolation Experiments
According to its technical reports [29], Qwen2.5-VL models are pre-trained with a context length of 32K tokens during the “Long-Context Pre-Training” stage. Meanwhile, its HuggingFace model card shows that we can use YaRN [37], with a scaling factor of 4, to extend its context length to 128K tokens [^7] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct#processing-long-texts. In Figure 14, we test the performance of using YaRN. We have two observations: (1) Using YaRN may hurt the performance on shorter input lengths. For example, at 8K tokens, the DocVQA score of Qwen2.5-VL-32B decreases from 67.8% to 63.4%; (2) On average across the entire MMLongBench, YaRN only substantially improves the performance of Qwen2.5-VL-3B (from 21.2 to 30.2 at 128K). However, the SoTA performance from large models (i.e., 32B and 72B) only fluctuates slightly.
To ensure a fair comparison, we do not apply YaRN to Qwen2.5-VL in our main evaluations. Since YaRN is used in a zero-shot way here, applying it would lead to an unfair comparison over other models that do not employ YaRN.
Figure 15: Results of applying V2PE [52] to InternVL2-2B. We find V2PE is very sensitive to the visual increment $\delta$ and overfitted to NIAH tasks. Note that we use ROUGE-L here. It is already sufficient to distinguish between models, so there is no need to use the costly GPT-4o evaluation. The numbers in parentheses (i.e., 16, 64, and 256) correspond to the visual increment $\delta\in\{\frac{1}{16},\frac{1}{64},\frac{1}{256}\}$, respectively.
@@
Ge et al. [52] proposed a positional embedding extrapolation method called V2PE, where it assigns smaller positional increments to visual tokens than textual tokens. They further applied V2PE to InternVL2 and trained the model to enhance its performance on MM-NIAH-Ret (I). In this experiment, we evaluate the V2PE-256K checkpoints [^8] https://huggingface.co/OpenGVLab/V2PE/tree/main/V2PE-256K with the visual increments $\delta\in\{\frac{1}{16},\frac{1}{64},\frac{1}{256}\}$. As shown in Figure 15, we find that (1) V2PE is very sensitive to different visual increments. For example, when we use $\frac{1}{16}$ and $\frac{1}{64}$, the performance is even worse than InternVL2-2B, which leads to extra hyperparameter tuning; (2) The V2PE (256) shows extremely high performance on NIAH tasks, which can be attributed to the fact that it was trained on MM-NIAH. The sharp performance difference on NIAH tasks versus other categories suggests that V2PE is strongly overfitted to the MM-NIAH dataset.
@@ D.6 Lost in the Middle
Existing works found that text-pure LLMs often struggle to recall needles in the middle of the input sequence, named lost in the middle [56]. On our benchmark, we extend the previous analysis to text-image long-context tasks with input length up to 128K tokens. We place the needle at six different evenly spaced depths in the context (i.e., $[0,0.2,0.4,0.6,0.8,1.0]$) and evaluate the LCVLMs’ ability to retrieve it. In our study, the needle may be a gold passage, an image, or a key sentence. We show the results in Figure 16 for ViQuAE, Figure 17 for VH-Single, Figure 18 for MM-NIAH-Ret (T), Figure 19 for MM-NIAH-Ret (I), Figure 20 for MM-NIAH-Reason (I).
We observe a similar phenomenon in many LCVLMs on the text-image long-context tasks. For example, the InternVL3-14B in Figure 18 and Ovis2-34B in Figure 19 both exhibit much better performance when the needle is at depths 00 and $1.0$. Furthermore, as we extend the context to longer lengths (e.g., 128K tokens), we observe cases where the model tends to favor either the very beginning or the very end of the context, but not both simultaneously. For example, as shown in Figure 16, InternVL3-8B prefers the very beginning of the context (depth 00) at 128K tokens, whereas Qwen2.5-VL-72B favors the very end (depth $1.0$).
D.7 Error Analaysis Details
We conducted two error analyses in Section 4.4. We provide more details of those two analyses in this section. First, for MMLongBench-Doc, we used PyMuPDF  [^9] https://pymupdf.readthedocs.io to extract the plain text from PDF-formatted documents. For ViQuAE, the entity names are already provided in the dataset, since it is constructed based on TriviaQA. The text-pure LLMs we used for Qwen2.5-VL models are Qwen2.5-7B-Instruct and Qwen2.5-32B-Instruct, which are instruction-tuned versions.
@@ Appendix E Prompts and Data Examples
We list a few data examples with prompts in Figures 26, 27, 28, 29, 30 and 31. For NIAH tasks, we provide examples of both VH and MM-NIAH, as their input formats are different.
@@ Appendix F Limitation
For limitation of evaluated models, while we already provide an extensive coverage of 46 frontier LCVLMs, there are still some models that we cannot cover due to token efficiency or integration challenges of codebases, as we discussed in Section B.1. We leave those works for future study. Meanwhile, the largest open-source models we evaluated are up to 72B in size (Qwen2-VL-72B and Qwen2.5-VL-72B). As we discussed in Appendix C, our computational resources are limited to 8 $\times$ A100 (80G) GPUs; thereby it is hard to deploy and evaluate larger models with over 100B parameters at the input length of 128K tokens, such as Llama4 [4].
@@ Appendix G Broader Impact
The long-context ability of LVLMs unlocked a large range of applications, including understanding documents with hundreds of pages and reasoning over dozens of web pages automatically. This ability can also help users to summarize a long document or revise a large-scale code repository. Meanwhile, there are also a large number of instruction-following scenarios grounded in complex visual contexts, such as long-term dialogue with humans or dialogue-based navigation for robots. Looking ahead, our MMLongBench will serve as a standard evaluation for the whole community to benchmark new LCVLMs and to stimulate the development of models with more efficient vision-language token encodings, more robust position-extrapolation schemes, and improved OCR, multi-modal retrieval, and reasoning capabilities.

### v2

@@ A.5 Long-Document VQA
First, we find that there is a proportion of noisy answer format annotations. For example, a list answer like [‘Top 10 File Categories Sorted By Disk Space’, ‘Last 12 Months Modified Disk Space History’] is annotated as being in string format. Conversely, answers in string format are also annotated in list format, such as [‘PRIVACY SCREEN OPTIONS’]. Therefore, our first step with these datasets is to correct the mislabeled answer formats and discard the instances for which the correct answer format cannot be recovered.
Second, both datasets rely heavily on LLMs, such as GPT-4o, to extract the answer from model predictions. This leads to high evaluation costs and poses challenges for large-scale evaluation, like 46 models in our work. Then, we take a closer examination of different formats of answers: (1) For integer and float answers, we find that numbers can be extracted with regular expressions; (2) For string answers, if the answer is short (less than 5 words), we find that model predictions are also short. Thus, we can directly use automatic metrics like ROUGE F1 without the need for answer extractions; (3) As a result, only long-form string answers require LLM-based extraction; Since long-form string answers ($>$ 5 words) constitute only a small proportion of these datasets, we simply discard those instances to enable scalable evaluation without relying on GPT-4o for answer extraction. We find that the retained short string answers are mostly entity names; (4) Note that for list answers, we evaluate each element in the list (i.e., integer, float, or string elements) separately (then take average), and the evaluation rule is determined for each element by its type.
After all the filtering, we find both datasets remain well-distributed as shown in Figure 9.
@@
We follow previous works [5, 17] and employ the same rule-based scoring method that applies different strategies depending on the format of the reference answer: (1) For String format answers, we initially use regular expressions to determine whether the answers require exact matching (e.g., telephone numbers, email addresses, website addresses, filenames, times, dates, etc.). If the answer needs an exact match, we perform a substring exact match (SubEM) with a score of 0 or 1. Otherwise, we follow previous works [82, 63, 97, 20] and calculate ROUGE F1 scores; (2) For integer answers, we perform an exact match comparison, and the score is either 0 or 1; (3) For float answers, we treat the model prediction and gold reference as the same if the relative error is less than 1%; (4) For list answers, we evaluate each element separately based on its answer type and take the average. Here, we follow LongDocURL to use the Greedy List Match: for each element in the reference list, we compute its score against all elements in the prediction list and greedily select the highest score as its matching score. This metric does not require the predicted list to follow the same element order as the reference list, thereby providing greater tolerance in evaluation.
@@
The input lengths of DocVQA tasks are also easy to control. If an example exceeds a given length $L$, we truncate the document evenly from both sides while preserving the answer pages. If the document cannot fill the length $L$, we alternately pad the left and right sides with randomly sampled negative documents until the required length is reached. Notably, we may also truncate a few pages of the last padding document as needed to control the length at the granularity of pages, instead of documents.
@@ A.6 Image Resizing and Statistics
As we discussed in Section 3.3, the number of tokens per image is determined by the image size in our benchmark. In MM-NIAH, we find that many images from the OBELICS dataset are unnecessarily large (up to 8000$\times$6000 pixels) and are not text-rich. Then, we resize those images’ longer edge to 1024 pixels while preserving their aspect ratio.
@@ B.1 LVLMs Beyond our Evaluation
DeepSeek-VL [103], DeepSeek-VL2 [48], and Long-Llava [53] are also excellent, but they present additional challenges. These models do not provide a standard API in the HuggingFace Transformers framework [104], which we used for inference. Therefore, these models cannot be loaded directly via Transformers, and we need to develop them based on their GitHub repositories. As a result, integrating these models requires a prohibitive amount of engineering effort to adapt their codebases, making it impractical within our current scope. We leave their integration for future work.
@@ Appendix C Experimental Setup
For all open-source models, we evaluate them on an 8$\times$×A100 (80GB) GPU server. We use the HuggingFace Transformers framework [104] to deploy models and generate outputs. Since all models are instruction-tuned, we apply the chat templates to all datasets. We load models in BF16 with FlashAttention2 [105] for faster inference. The largest open-source models tested in our work have 72B parameters. Our computational resources are limited to 8$\times$×A100 GPUs; thereby, we cannot evaluate models with over 100B parameters, such as Llama4 [4], at 128K tokens.
We sampled 100 examples from each dataset to evaluate models. Note that in MM-NIAH, we sample the text-needle and image-needle examples evenly, with 50 of each type. This amount actually results in 600 examples for single-needle tasks: ViQuAE, VH-Single, and MM-NIAH-Ret, and 300 examples for multi-needle tasks: InfoSeek, VH-Multi, and MM-NIAH-Count. This is because we test 6 different depths (i.e., [0, 0.2, 0.4, 0.6, 0.8, 1.0]) for single-needle examples and 3 different permutations for multi-needle examples to mitigate the positional bias. The MM-NIAH-Reason is more complex since image-needle (I) examples have a single needle, while text-needle (T) ones have multiple needles. There are 300 examples for MM-NIAH-Reason (I) (50 $\times$ 6 depths) and 150 examples for MM-NIAH-Reason (T) (50 $\times$ 3 permutations). Due to this modality imbalance in MM-NIAH-Reason, we compute the average score for the subset of each modality separately and report their mean as the final result. Together, we evaluate each model on 4,050 examples across five different input lengths, resulting in a total of 20,250 examples.
Appendix D Additional Results
@@ D.1 The Difficulty of Visual Haystack
We manually checked a number of examples from the Visual Haystack dataset and didn’t find any errors in the task labels. As shown in Figure 10, Gemini-2.5-Pro achieves an accuracy of 85.4 on VH-Single, and GPT-4o achieves an accuracy of 70.3 on VH-Multi. These results are higher than a random guess (50%), which demonstrates the correctness of our implementation.
@@
D.2 Correlation between NIAH and Various Downstream Tasks
@@
We further show the performance of the selected models on the text-needle and image-needle subsets in Figure 12. We observe that models exhibit weak performance on MM-NIAH-Count (I) and MM-NIAH-Reason (I). This challenging nature leads to a low degree of separability between different models.
@@ D.4 Performance of Claude
At the time of our evaluation, the Claude 3 family of models can take up to 100 images [^6] https://docs.anthropic.com/en/docs/build-with-claude/vision#basics-and-limits in a single request. However, a few datasets, such as Food101, VH-Single, or SlildeVQA, contain hundreds of images at input lengths of 64K and 128K tokens. As a result, it is impossible to process all images in a single pass through the model. For each input length, if one or more datasets within a category contain samples with more than 100 images, we exclude that category from evaluation at that input length. We provide the statistics about the average image number per example in each dataset at all five input lengths in Table 3 and Section A.6. When testing Claude-3.7-Sonnet, we mark those untestable cases as “N/A” in our results to distinguish them from genuine model failures (which receive a score of 0).
Figure 14: Results of applying YaRN [37] to Qwen2.5-VL models. We find that YaRN only improves the performance of Qwen2.5-VL-3B substantially. However, the SoTA performance from larger models (i.e., 32B and 72B) only fluctuates slightly.
@@ D.5 Positional Embedding Extrapolation Experiments
According to its technical reports [29], Qwen2.5-VL models are pre-trained with a context length of 32K tokens during the “Long-Context Pre-Training” stage. Meanwhile, its HuggingFace model card shows that we can use YaRN [37], with a scaling factor of 4, to extend its context length to 128K tokens [^7] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct#processing-long-texts. Note that YaRN is used in a zero-shot manner, as there is no continual training for YaRN on Qwen2.5-VL. In Figure 14, we test the performance of using YaRN. We have two observations: (1) Using YaRN may hurt the performance on shorter input lengths. For example, at 8K tokens, the DocVQA score of Qwen2.5-VL-32B decreases from 67.8% to 63.4%; (2) On average across the entire MMLongBench, YaRN only substantially improves the performance of Qwen2.5-VL-3B (from 21.2 to 30.2 at 128K). However, the SoTA performance from large models (i.e., 32B and 72B) only fluctuates slightly.
To ensure a fair comparison, we do not apply YaRN to Qwen2.5-VL in our main evaluations. Since YaRN is used in a zero-shot way here, applying it would lead to an unfair comparison over other models.
Figure 15: Results of applying V2PE [52] to InternVL2-2B. We find V2PE is very sensitive to the visual increment $\delta$ and overfitted to NIAH tasks. Note that we use ROUGE-L for summarization here, since it is already sufficient to distinguish between models. Thus, there is no need to use the costly GPT-4o evaluation. The numbers in parentheses (i.e., 16, 64, and 256) correspond to the visual increment $\delta\in\{\frac{1}{16},\frac{1}{64},\frac{1}{256}\}$, respectively.
@@
Ge et al. [52] proposed a positional embedding extrapolation method called V2PE, where it assigns smaller positional increments to visual tokens than textual tokens. They further applied V2PE to InternVL2-2B and trained the model to enhance its performance on MM-NIAH-Ret (I). In this experiment, we evaluate the V2PE-256K checkpoints [^8] https://huggingface.co/OpenGVLab/V2PE/tree/main/V2PE-256K with the visual increments $\delta\in\{\frac{1}{16},\frac{1}{64},\frac{1}{256}\}$. As shown in Figure 15, we find that (1) V2PE is very sensitive to different visual increments. For example, when we use $\frac{1}{16}$ and $\frac{1}{64}$, the performance is even worse than InternVL2-2B, which leads to extra hyperparameter tuning; (2) The V2PE (256) shows extremely high performance on NIAH tasks, which can be attributed to the fact that it was trained on MM-NIAH. The sharp performance difference on NIAH tasks versus other categories suggests that V2PE is strongly overfitted to the MM-NIAH dataset.
@@ D.6 Lost in the Middle
Existing works found that text-pure LLMs often struggle to recall needles in the middle of the input sequence, named lost in the middle [57]. On our benchmark, we extend the previous analysis to vision-language tasks with input length up to 128K tokens. We place the needle at six different evenly spaced depths in the context (i.e., $[0,0.2,0.4,0.6,0.8,1.0]$) and evaluate the LCVLMs’ ability to retrieve it. In our study, the needle may be a gold passage, an image, or a key sentence. We show the results in Figure 16 for ViQuAE, Figure 17 for VH-Single, Figure 18 for MM-NIAH-Ret (T), Figure 19 for MM-NIAH-Ret (I), Figure 20 for MM-NIAH-Reason (I).
We observe a similar lost-in-the-middle phenomenon in many LCVLMs on long-context vision-language tasks. For example, the InternVL3-14B in Figure 18 and Ovis2-34B in Figure 19 both exhibit much better performance when the needle is at depths 00 and $1.0$. Furthermore, as we extend the context to longer lengths (e.g., 128K tokens), we observe cases where the model tends to favor either the very beginning or the very end of the context, but not both simultaneously. For example, as shown in Figure 16, InternVL3-8B prefers the very beginning of the context (depth 00) at 128K tokens, whereas Qwen2.5-VL-72B favors the very end (depth $1.0$).
D.7 Error Analysis Details
We conducted two error analyses in Section 4.4. We provide more details of those two analyses in this section. First, for MMLongBench-Doc, we used PyMuPDF  [^9] https://pymupdf.readthedocs.io to extract the plain text from PDF-formatted documents. For ViQuAE, the entity names are already provided in the dataset, since it is constructed based on TriviaQA. The text-pure LLMs we used, corresponding to Qwen2.5-VL models, are Qwen2.5-7B-Instruct and Qwen2.5-32B-Instruct, which are instruction-tuned versions.
@@ Appendix E Prompts and Data Examples
We list a few data examples with prompts in Figures 26, 27, 28, 29, 30 and 31. For NIAH tasks, we provide examples of both VH and MM-NIAH, as their input formats are very different.
@@ Appendix F Limitation
For limitation of evaluated models, while we already provide an extensive coverage of 46 frontier LCVLMs, there are still some models that we cannot cover due to token efficiency or integration challenges of codebases, as we discussed in Section B.1. We leave those works for future study. Meanwhile, the largest open-source models we evaluated are up to 72B in size (Qwen2-VL-72B and Qwen2.5-VL-72B). As we discussed in Appendix C, our computational resources are limited to 8$\times$×A100 (80G) GPUs; thereby it is hard to deploy and evaluate larger models with over 100B parameters at the input length of 128K tokens, such as Llama4 [4].
@@ Appendix G Broader Impact
The long-context ability of LVLMs has unlocked a large range of applications, including understanding documents with hundreds of pages and reasoning over dozens of web pages automatically. This ability can also help users to summarize a long document or revise a large-scale code repository. Meanwhile, there are a large number of instruction-following scenarios grounded in complex vision-language contexts, such as long-term dialogue with humans or dialogue-based navigation for robots. Looking ahead, our MMLongBench will serve as a standard evaluation for the whole community to benchmark new LCVLMs and to stimulate the development of models with more efficient vision-language token encodings, more robust position-extrapolation schemes, and improved OCR, multi-modal retrieval, and reasoning capabilities.